OpenStudy (anonymous):

Q: Suppose are driving to visit a friend in another state. You are driving at an average rate of 50 mile per hour. You must drive a total of 345 miles. If you have already driven 145 miles, how long will it take you to reach your destination? A: 5 Hours B: 3 Hours C: 4 Hours D: 6 Hours

OpenStudy (anonymous):

help plz

OpenStudy (bibby):

You have to drive 345-145 miles at 50 miles per hour. Does that simplify the question?

OpenStudy (anonymous):

so 345 -145 * 50

OpenStudy (bibby):

that would not match the answers in either units or numbers., gimme a bit, playin a game

OpenStudy (bibby):

Someone else will pickup I s'pose

OpenStudy (nurali):

Applying "distance formula": d = rt where d is distance r is rate or speed t is time Solving for 't': t = d/r t = (345-145)/50 t = 200/50 t = 4 hours